There are four green blocks and four red blocks in a bag.
a What is the probability of selecting a red block, keeping it, then selecting
another red block?
b What is the probability of selecting a green block, replacing it, then selecting a red block?

Answer by stanbon(57214)

There are four green blocks and four red blocks in a bag.
a What is the probability of selecting a red block, keeping it, then selecting
another red block?
---
1st red: 4/8
2nd red: 3/7
---
P(red/red) = (4/8)*(3/7) = 3/14
-----------------------------------------
b What is the probability of selecting a green block, replacing it, then selecting a red block?
---
P(green/red) = (1/2)(1/2) = 1/4
===================================
Cheers,
Stan H.
================